a piece of wire is 0.524m long. it is placed at right angles to form a uniform magnetic field o 0.25T. The force on the wire is 0.063N. calculate the current in the wire
1
Expert's answer
2018-02-21T04:15:07-0500
According to Ampère's force law (for the straight wire)
F=IBL sin⁡α,
we get
I=F/(BL sin⁡α )=0.063/(0.25∙0.524∙sin⁡〖90°〗 )=0.481 A.